1. (x+3y-5)dx-(x-y-1)dy=0


1
Expert's answer
2021-01-05T17:27:12-0500

Substitutions X=x-2, Y=y-1 is clear from solving stationary point equations:

"\begin{aligned}
 x+3y-5=0 \\
 x-y-1=0
\end{aligned}" => "\begin{aligned}
 x=2 \\
 y=1
\end{aligned}"

Equation in terms of the variable z was integrated as follows:


"\frac{dz}{dX}X=\frac{(z+1)^2}{1-z}" => "\frac{(1-z)dz}{(z+1)^2}=\frac{dX}{X}" => "\frac{2-(1+z)}{(z+1)^2}dz=\frac{dX}{X}"

Integrating

"\int\frac{2-(1+z)}{(z+1)^2}dz=\int\frac{dX}{X}" => "\int\frac{2dz}{(z+1)^2}-\int\frac{dz}{z+1}=\int\frac{dX}{X}" =>"-\frac{2}{z+1}-ln|z+1|=ln|X|-C" => "C=\frac{2}{z+1}+ln|z+1|+ln|X|"

Remember z=Y/X. So from here

"C=\frac{2}{\frac{Y}{X}+1}+ln|\frac{Y}{X}+1|+ln|X|" => "C=\frac{2X}{X+Y}+ln|X+Y|"